OpenStudy (anonymous):

4.2 x 10^-4 in standard form

OpenStudy (amorfide):

since the power of 10 is negative, you know that you are actually dividing 4.2 by this number therefore it gets smaller \[a^{-n}=\frac{ 1 }{ a^n }\]

OpenStudy (amorfide):

so you are dividing by 10, 4 times

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

im doing scientific notation @amorfide

OpenStudy (misty1212):

take the number \(4.2\) and move the decimal point 4 places to the left

OpenStudy (amorfide):

it is already in scientific form

OpenStudy (amorfide):

you want standard form, which is its actual number

OpenStudy (anonymous):

yea so i need it back in standard form... so 40? @amorfide

OpenStudy (misty1212):

oh no!\[0.\overleftarrow{ 0004}\]

OpenStudy (misty1212):

move the decimal point four places to the LEFT

OpenStudy (misty1212):

it was here \[\huge 4\color{red}.2\] now it is here \[\huge 0\color{red}.\overleftarrow{ 0004}2\]

OpenStudy (amorfide):

\[10^{-4}=\frac{ 1 }{ 10 } \times \frac{ 1 }{ 10 } \times \frac{ 1 }{ 10 } \times \frac{ 1 }{ 10 } \] so you have \[4.2 \times 10^{-4}= 4.2 \times \frac{ 1 }{ 10 } \times \frac{ 1 }{ 10 } \times \frac{ 1 }{ 10 } \times \frac{ 1 }{ 10 }\]

OpenStudy (misty1212):

\[\huge 0\color{red}.\overleftarrow{\overbrace{ 0004}}2\]

OpenStudy (anonymous):

im sorry bit im just confused...@misty1212

OpenStudy (misty1212):

i wish i knew another way to say it the exponent is \(-4\) take the decimal point on move it four places to the left that is all

OpenStudy (anonymous):

42000 or .420000

OpenStudy (amorfide):

okay so let us say that the power is denoted by x x=-4 -4 on the number line is to the left of zero therefore since your power of 10 is negative 4 you must move the decimal point 4 places to the left you have 4.2 x 10^-4 so 4.2 moved one place to the left is 0.42 0.42 one place to the left is 0.042 0.042 one place to the left is 0.0042 0.0042 one place to the left is 0.00042